# SAGE: Sequence-level Adaptive Gradient Evolution for Generative Recommendation

### v1

@@ 3.3.2. Multi-Objective Signal Decoupling
Decoupled Slate Advantage Directly summing weighted rewards leads to ”Reward Collapse,” where distinct behaviors map to identical advantage values. Instead, we normalize each objective independently within the sampling group before aggregation:
@@
  • $R_{m}(S)$: The raw reward value for the $m$-th objective (e.g., watch time).
  • $\mu_{m}^{group},\sigma_{m}^{group}$: The mean and standard deviation of rewards within the sampled group $G$ for a single user query.
  • $\text{Norm}_{batch}$: A final batch-level normalization ($z$-score) to ensure $A_{SAGE}$ follows a standard normal distribution, providing a stable scale for the gradient boundaries.
Sequence-level Ratio Formulation Since user feedback is directed at the item level, token-level importance ratios introduce high variance. We define the importance ratio as the Geometric Mean of the token probability ratios over the generated list:
@@
  • Physical Meaning: This metric smooths out intra-item token fluctuations, representing the average probability gain of generating the entire slate under the new policy versus the old policy.
4. Experiments
In this section, we conduct comprehensive experiments to evaluate the effectiveness of the proposed framework. We aim to validate whether our Native-LLM + SAGE paradigm (”RecLLM”) can outperform state-of-the-art generative baselines in terms of accuracy, while simultaneously resolving the cold-start and diversity bottlenecks inherent in the OneRec-V2 paradigm.
4.1. Experimental Setup
4.1.1. Datasets
Public Datasets: We evaluated the proposed framework on three publicly available real-world benchmarks from the Amazon Product Reviews dataset  (He and McAuley, 2016), which encompasses user reviews and item metadata spanning from May 1996 to July 2014. Specifically, we utilized three categories of this dataset for the sequential recommendation task: “Sports” “Beauty” and “Toys”.
4.1.2. Baselines
We compare our method against distinct categories of baselines:
  • SASRec: The classic self-attentive sequential model.
  • Generative/Retrieval Models: ReaRec, TIGER, and LC-Rec, which utilize VQ-VAE or retrieval-augmented generation.
  • OneRec Variants:
    • OneRec-V2: The SOTA generative baseline using Semantic IDs and GBPO optimizer.
    • OneRec-SAGE: Applying our SAGE optimizer to the traditional OneRec (Semantic ID) architecture to test optimizer effectiveness in isolation.
    • Only SFT (RecLLM): The Native-LLM model trained only with Supervised Fine-Tuning.
    • RecLLM: The proposed framework combining the Native-LLM architecture with the SAGE optimizer.
4.1.3. Evaluation Metrics
We evaluate performance across three dimensions using top-$K$ metrics (default $K=10$):
  • Accuracy: Normalized Discounted Cumulative Gain (NDCG@K) and Recall (Recall@K).
  • Diversity: Intra-List Diversity (ILD), measuring the average semantic distance between items within a generated slate, and Entropy@K, measuring the distribution of recommended categories across the test set.Since all items in a dataset category belong to the same root (e.g., ”Beauty”), calculating entropy at the root level is trivial. Therefore, we utilize the second-level sub-categories (e.g., ”Skin Care”, ”Makeup”, ”Fragrance” for Beauty) provided in the item metadata to calculate Shannon Entropy. This measures whether the model recommends a diverse range of granular interests rather than converging to a single sub-niche.
  • Cold-Start: Given that precise item publication timestamps are unavailable in this dataset, we adopt the standard academic definition for cold-start evaluation. We define Cold Items as the set of items belonging to the bottom 20% of the interaction frequency distribution in the training set. This metric evaluates the model’s ability to discover and recommend long-tail items that have sparse historical interactions, preventing the ”rich-get-richer” feedback loop.
4.1.4. Implementation Details
The base model is Qwen3-8B. We set the slate length $L=6$. The SFT stage is trained for 1 epochs. In the RLHF stage, we use a learning rate of 5e65e-6 with the AdamW optimizer. For SAGE specific hyperparameters, based on validation set tuning, we set the Positive Boost Factor $\epsilon_{boost}=0.3$, the Diversity Temperature $\beta=0.5$.
4.2. Main Results
Superiority of RecLLM: As shown in Table  1, our proposed method (”RecLLM”) consistently outperforms all baselines across all datasets. Specifically, compared to the strongest baseline OneRec-V2, RecLLM achieves improvements of 4.3% to 8.9% in NDCG metrics. This confirms that directly reusing the Native LLM vocabulary—when optimized correctly—can surpass models relying on heavy, separate tokenization.
The Necessity of SAGE: Notably, the ”Only SFT” model performs poorly, significantly lagging behind even SASRec. SFT mainly functions to activate the model’s fundamental capabilities. without reinforcement learning to align the Native LLM with recommendation goals, the model fails to capture complex user preferences. Furthermore, OneRec-SAGE (applying SAGE to the old architecture) shows mixed results: while competitive, it slightly underperforms OneRec-V2 in some metrics (e.g., R@5 on Beauty). This suggests that SAGE’s aggressive exploration requires the robust semantic representation of the RecLLM to fully realize its potential without destabilizing the training.
Table  2 investigates the impact of optimization strategies on recommendation diversity (Entropy) and long-tail item discovery (Cold-Recall).
Breaking the Symmetric Conservatism: Comparing RecLLM against OneRec-V2, we observe substantial gains:
  • Cold-Recall: SAGE improves cold-start performance by 49% (Beauty), 63% (Sports) and 43% (Toys). This validates that the Positive Boost Factor ($\epsilon_{boost}$) effectively allows high-potential fresh items to break through the symmetric gradient barrier of GBPO.
  • Diversity: Entropy improves by 5.9% , 10.36% and 7.36%, confirming that the Entropy-Aware Penalty prevents the model from collapsing into safe, popular items.
The Accuracy-Diversity Trade-off: It is worth noting that OneRec-SAGE achieves the highest absolute values for Entropy and Cold-Recall. However, revisiting Table  1, we see that OneRec-SAGE suffers a drop in top-k accuracy (e.g., lower R@5 than OneRec-V2). This indicates that applying aggressive exploration on the limited semantic ID vocabulary can lead to over-exploration. In contrast, RecLLM achieves the most Optimal state: it delivers the highest overall accuracy table 1 while still providing a massive upgrade in diversity and cold-start handling compared to the current SOTA baseline (OneRec-V2).
@@ 4.2. Main Results
OneRec-V2 0.0646 0.0929 0.0456 0.0550 0.0365 0.0551 0.0252 0.0317 0.0693 0.0958 0.0496 0.0581
OneRec-SAGE 0.0612 0.0931 0.0424 0.0557 0.0351 0.0541 0.0246 0.0297 0.0671 0.0960 0.0466 0.0583
Only SFT (RecLLM) 0.0291 0.023 0.0158 0.204 0.0163 0.0263 0.0104 0.0101 0.0386 0.0571 0.0169 0.0212
RecLLM 0.0683 0.0973 0.0497 0.0590 0.0385 0.0575 0.0271 0.0339 0.0738 0.1018 0.0533 0.0622
Improve +5.67% +4.7% +8.9% +7.2% +5.46% +4.33% +7.35% +6.96% +6.5% +6.3% +7.48% +7.11%
Table 2. Relative to OneRec‑v2, SAGE delivers significant gains in cold‑start recovery and recommendation diversity.
@@
Method Entropy@10 Cold-Recall Entropy@10 Cold-Recall Entropy@10 Cold-Recall
OneRec-V2 2.285 0.0326 2.274 0.0282 2.315 0.0365
OneRec-SAGE 2.551 0.0619 2.528 0.0713 2.574 0.0695
RecLLM 2.420 0.0487 2.510 0.0460 2.485 0.0522
Improve +5.9% +49% +10.36% +63% +7.36% +43%
@@ 4.3. Ablation Study
To understand the contribution of each component, we conduct an ablation study on the Amazon-Beauty dataset. The results are visually presented in Figure  5.
  • w/o Positive Boost: Setting $\epsilon_{boost}=0$ reverts the positive boundary to standard GBPO ($r\leq 1$). We observe a drastic drop in Cold-Start Recall (Green Bar) to 0.50 of the full model’s performance, confirming this component’s critical role in item exploration.
  • w/o Entropy Penalty: Removing the entropy term and using the static GBPO boundary for negative feedback leads to a significant decrease in Entropy@10 (Orange Bar) to 0.78, indicating a collapse towards redundant recommendations.
  • w/o Signal Decoupling: Using a simple weighted sum of rewards instead of decoupled normalization results in a noticeable drop in NDCG@10 (Blue Bar) to 0.90 and a decrease in other metrics. This suggests that ”Reward Collapse” was indeed hindering the precise multi-objective optimization.
The full SAGE model achieves the best balance, demonstrating that the asymmetric adaptive dynamics and signal decoupling are complementary and essential.
@@ 5. Conclusion
In this paper, we identified fundamental limitations in existing optimization strategies for generative recommendation, specifically the ”Symmetric Conservatism” of Gradient-Bounded Policy Optimization (GBPO), which leads to cold-start suppression and diversity collapse. To address these challenges within a list-wise generation paradigm, we proposed SAGE (Sequence-level Adaptive Gradient Evolution).
SAGE introduces a novel theoretical framework featuring **Asymmetric Adaptive Dynamics**. By applying a ”Positive Boost” to high-potential successful slates and an ”Entropy-Aware Penalty” to low-diversity failures, SAGE dynamically aligns the optimization trajectory with the long-tailed, high-noise nature of recommendation feedback. Furthermore, we incorporated Multi-Objective Signal Decoupling to ensure high-fidelity gradient signals in complex real-world scenarios.
RecLLM demonstrates that native‑vocabulary reuse not only surpasses models dependent on large standalone tokenizers but also maximizes the utility of the underlying LLM. While the absence of an independent vocabulary necessitates language‑level item encoding—posing challenges for real‑time systems—this constraint is expected to fade with increasing computational capacity.

### v2

@@ 3.4.3. Multi-Objective Signal Decoupling
Decoupled Slate Advantage. Directly summing weighted rewards can cause Reward Collapse, where distinct behaviors map to identical advantage values. To avoid this, we normalize each objective independently within the sampling group and then aggregate them. For a slate $S$, the advantage is defined as
@@
where $R_{m}(S)$ denotes the raw reward of the $m$-th objective (e.g., Click or Comment), and $\mu_{m}^{group},\sigma_{m}^{group}$ are the mean and standard deviation computed within the sampled group $G$ for the same user query. The operator $\text{Norm}_{batch}$ applies a final batch-level $z$-score normalization so that $A_{SAGE}$ stays on a stable scale (approximately standard normal), which helps maintain well-behaved gradient magnitudes.
Sequence-level Ratio Formulation. Since user feedback is directed at the item level, token-level importance ratios can introduce high variance. We therefore define the importance ratio at the slate level as the geometric mean of token probability ratios across the generated list:
@@
with a fixed slate length $L=10$. Physical Meaning: this formulation smooths intra-item token fluctuations and can be interpreted as the average probability gain of generating the entire slate under the new policy relative to the old policy.
@@
OneRec-GRPO 0.0646 0.0924 0.0456 0.0545 0.0365 0.0547 0.0252 0.0310 0.0693 0.0953 0.0496 0.0579
OneRec-GBPO 0.0649 0.0928 0.0461 0.0550 0.0368 0.0551 0.0255 0.0317 0.0698 0.0958 0.0499 0.0581
OneRec-SAGE 0.0651 0.0932 0.0463 0.0557 0.0371 0.0553 0.0256 0.0320 0.0701 0.0960 0.0502 0.0585
Only SFT (TextRec) 0.0291 0.0495 0.0158 0.0204 0.0163 0.0263 0.0104 0.0139 0.0386 0.0571 0.0169 0.0212
TextRec 0.0683 0.0973 0.0497 0.0590 0.0385 0.0575 0.0271 0.0339 0.0738 0.1018 0.0533 0.0622
Improve +5.23% +4.84% +7.8% +7.27% +4.62% +4.35% +6.27% +6.94% +5.73% +6.26% +6.81% +7.05%
Table 2. Relative to OneRec‑GBPO, SAGE delivers significant gains in cold‑start recovery and recommendation diversity.
@@
Method Entropy@10 Cold-Recall@10 Entropy@10 Cold-Recall@10 Entropy@10 Cold-Recall@10
OneRec-GBPO 2.285 0.326 2.274 0.282 2.315 0.365
OneRec-SAGE 2.551 0.619 2.528 0.567 2.574 0.705
TextRec 2.420 0.487 2.510 0.460 2.485 0.522
Improve +11.64% +89.9% +11.17% +101% +11.19% +93.2%
Table 3. Performance on RecIF-Bench. Bold indicates the best and underline indicates the second best.
Method Short Video Rec Ad Rec Product Rec Label Pred. Interactive Rec Label-Cond. Rec
Pass@1 Pass@32 Recall@32 Pass@1 Pass@32 Recall@32 Pass@1 Pass@32 Recall@32 AUC Pass@1 Pass@32 Recall@32 Pass@32 Recall@32
SASRec 0.0045 0.1003 0.0119 0.0044 0.0980 0.0293 0.0052 0.0914 0.0175 0.6244 - - - 0.0380 0.0140
ReaRec 0.0052 0.1002 0.0120 0.0035 0.1054 0.0327 0.0030 0.0907 0.0189 0.6204 - - - 0.0381 0.0137
TIGER 0.0168 0.1061 0.0132 0.0125 0.1769 0.0581 0.0120 0.1276 0.0283 0.6675 - - - 0.0337 0.0123
LC-Rec 0.0341 0.1306 0.0180 0.0197 0.2096 0.0723 0.0178 0.1809 0.0416 0.6139 0.0890 0.3730 0.2394 0.0420 0.0170
OneRec-GRPO 0.0548 0.2122 0.0369 0.0259 0.2700 0.0964 0.0223 0.2290 0.0538 0.6912 0.1250 0.5080 0.3458 0.0549 0.0235
OneRec-GBPO 0.0551 0.2134 0.0377 0.0264 0.2731 0.0979 0.0227 0.2298 0.0545 0.6918 0.1256 0.5096 0.3481 0.0553 0.0248
OneRec-SAGE 0.0554 0.2142 0.0381 0.0266 0.2733 0.0983 0.0229 0.2302 0.0554 0.6919 0.1258 0.5107 0.3495 0.0561 0.0254
Only SFT (TextRec) 0.0049 0.0920 0.0107 0.0041 0.0940 0.0235 0.0032 0.0959 0.0161 0.6028 0.0410 0.2680 0.1660 0.0334 0.0128
TextRec 0.0574 0.2196 0.0392 0.0273 0.2798 0.1015 0.0231 0.2410 0.0568 0.7017 0.1310 0.5280 0.3648 0.0574 0.0261
4. Experiments
We conduct experiments on both public academic benchmarks and a large-scale industrial benchmark to evaluate SAGE as a general-purpose optimizer for list-wise generative recommendation. Our goal is to verify that SAGE (i) improves accuracy while unlocking cold-start learning signals, (ii) mitigates diversity collapse under rejection-dominated feedback, and (iii) remains numerically stable. We evaluate SAGE across two representative action spaces: Semantic-ID generation (OneRec-style) and native-text generation (TextRec-style).
4.1. Experimental Setup
4.1.1. Datasets
Amazon Product Reviews. We evaluate sequential recommendation on three real-world subsets from the Amazon Product Reviews dataset  (McAuley et al., 2015), which contains user reviews and item metadata from May 1996 to July 2014. We use the three categories Sports, Beauty, and Toys.
Table 4. Diversity and cold-start evaluation on RecIF-Bench.
Method Short Video
Entropy@10 Cold-Recall@10
OneRec-GBPO 4.284 0.521
OneRec-SAGE 4.455 0.637
TextRec 4.317 0.594
RecIF-Bench. In addition, we evaluate on RecIF-Bench  (Zhou et al., 2025a), an instruction-following benchmark for generative recommendation foundation models. RecIF-Bench contains approximately 120M interactions from 200K users across three industrial domains: Short Video (impression sequences with multi-behavior feedback such as like/follow/comment/effective view/dislike), Ad (click sequences), and Product (click sequences). The benchmark provides (i) user portraits written as narrative text interleaved with behavioral evidence, (ii) item-side metadata and multimodal representations (text and visual embeddings; dense captions for a large portion of videos), and (iii) interaction-side labels enabling behavior-conditional modeling.
We follow the official strict user-based split protocol: 20% users are held out for testing with zero user leakage; for each user, the interaction sequence is temporally split into history context and future targets. We report results on the core five recommendation tasks (Short Video Rec, Ad Rec, Product Rec, Interactive Rec, and Label-Conditional Rec) and the pointwise Label Prediction task.
Native-text action space (selection within candidate pool).
In TextRec, the policy does not freely generate arbitrary item strings. Instead, for each context $c$ we provide a candidate pool $\mathcal{I}(c)$ with textual item renderings (e.g., title/tags/category keywords), and the model outputs a length-$L$ slate by selecting items from $\mathcal{I}(c)$ (with deterministic parsing and validity checks). Outputs that do not match any candidate are treated as invalid and discarded (or mapped to a special null action), ensuring a well-defined and reproducible action space.
4.1.2. Baselines
To validate the effectiveness of our approach, we compare against three groups of competitive baselines, covering classical discriminative sequential recommenders, recent generative/retrieval-augmented recommenders, and OneRec-family variants.
Discriminative sequential recommenders: SASRec (Kang and McAuley, 2018). These methods are inherently task-specific, and thus require training a separate model for each RecIF task.
Generative / retrieval-augmented recommenders: ReaRec (Tang et al., 2025), TIGER (Rajput et al., 2023), and LC-Rec (Zheng et al., 2024). Following prior practice, we adapt these methods with task-specific modifications to support different RecIF tasks. For a fair comparison at the foundation-model scale, we implement LC-Rec as LC-Rec-8B by instantiating it with a comparable Qwen3-8B backbone.
OneRec variants and our method: OneRec-GRPO (Zhou et al., 2025a) (Semantic IDs + GRPO), OneRec-GBPO (Zhou et al., 2025b) (Semantic IDs + GBPO), OneRec-SAGE (applying SAGE on the Semantic-ID architecture), Only SFT (TextRec) (native-vocabulary model trained only with supervised fine-tuning), and TextRec (native-vocabulary model further optimized with SAGE).
4.1.3. Evaluation Metrics
Amazon metrics. We evaluate top-$K$ performance (default $K=10$) from three perspectives: (i) Accuracy using NDCG@K and Recall@K; (ii) Diversity using Entropy@K. Since all items in a dataset category share the same root label (e.g., ”Beauty”), entropy at the root level is trivial; therefore, we compute Shannon Entropy using second-level sub-categories from item metadata (e.g., Skin Care, Makeup, Fragrance); (iii) Cold-start using Cold-Recall@K, where cold items are defined as the bottom 5% of the interaction-frequency distribution in the training set.
RecIF-Bench metrics. We follow the official RecIF-Bench evaluation  (Zhou et al., 2025a). For recommendation tasks, we using Pass@K and Recall@K with $K\in\{1,32\}$. For Label Prediction, we using AUC.
4.1.4. Implementation Details
Backbone and generation. The base model is Qwen3-8B. We set slate length to $L=10$.
Training recipe. We use a two-stage pipeline (SFT then RLHF) as described in Section 3. The SFT stage is trained for 1 epoch. In the RLHF stage, we use AdamW with a learning rate of 5e65e-6. For SAGE hyperparameters, based on validation tuning, we set the Positive Boost Factor $\epsilon_{boost}=0.3$ and Diversity Temperature $\beta=0.5$.
RecIF-Bench specific settings. For TextRec, items are represented purely by native-text (title/tags/category keywords) without introducing a separate item-token vocabulary. For interactive and conditional recommendation, $L=10$, and we generate candidates using nucleus sampling with temperature 0.9. For OneRec-GRPO, OneRec-GBPO and OneRec-SAGE, we use the public checkpoints and evaluation settings in OpenOneRec  (Zhou et al., 2025a). For task-specific sequential baselines and TIGER, we reuse the reported benchmark numbers under the same protocol.
4.2. Main Results
4.2.1. Results on Amazon Product Reviews
Table 1 shows that TextRec consistently achieves the best top-$K$ accuracy across Beauty, Sports, and Toys. Compared to the strongest baseline OneRec-GBPO, TextRec improves NDCG by 6.27% to 7.8%, indicating that directly reusing the native LLM vocabulary—when paired with preference optimization—can outperform architectures relying on separate Semantic-ID tokenizers.
The Only SFT (TextRec) variant performs poorly and lags behind SASRec, suggesting that SFT mainly teaches formatting and coarse preference patterns; without preference optimization, the model fails to align generation with ranking objectives. Applying SAGE to the Semantic-ID architecture (OneRec-SAGE) yields competitive but mixed accuracy changes, implying that aggressive exploration is harder to stabilize when the output space is constrained by a discrete Semantic-ID vocabulary.
Beyond accuracy, Table 2 highlights SAGE’s effect on long-tail discovery and diversity. Relative to OneRec-GBPO, SAGE yields large gains in Cold-Recall and also improves Entropy, supporting our diagnosis that GBPO-style symmetric clipping tends to suppress positive learning signals for rare items and encourages conservative, popularity-biased updates.
4.2.2. Results on RecIF-Bench
Table 3 reports results on six RecIF-Bench tasks. TextRec consistently outperforms OneRec-GBPO across the core recommendation tasks, with improvements that are especially clear on instruction-conditioned settings (e.g., Interactive Rec and Label-Conditional Rec). This indicates that representing items with native-text can better exploit the backbone LLM’s semantic priors when the task is framed as instruction following rather than pure next-item prediction.
The Only SFT variant again underperforms, consistent with the Amazon findings: SFT alone does not provide sufficiently strong preference signals to drive high-quality ranking under the RecIF-Bench protocol. OneRec-SAGE remains competitive but does not consistently dominate top-$K$ accuracy, reinforcing the empirical pattern that exploration can become over-aggressive when the action space is limited by a Semantic-ID vocabulary.
Finally, Table 4 evaluates diversity and cold-start on RecIF-Bench (Short Video). Relative to OneRec-GBPO, TextRec improves both Entropy@10 and Cold-Recall@10, while OneRec-SAGE achieves the highest absolute values on these two metrics. Taken together with the accuracy table, this suggests a consistent trade-off: SAGE increases exploration and long-tail coverage; the native-vocabulary TextRec setting better preserves accuracy under such exploration than Semantic-ID generation.
@@ 4.3. Ablation Study
To isolate the contribution of each component, we conduct ablations on Amazon-Beauty. Figure 5 normalizes metrics to the full model.
Removing Positive Boost ($\epsilon_{boost}=0$) drastically reduces Cold-Start Recall, showing that relaxing the conservative positive update boundary is crucial for long-tail discovery. Removing the Entropy Penalty substantially decreases Entropy@10, indicating that diversity-aware negative feedback prevents collapse into redundant popular outputs. Replacing Signal Decoupling with a simple weighted reward sum reduces NDCG@10 and other metrics, supporting the hypothesis that multi-objective reward mixing can cause signal interference and unstable optimization. Overall, the best results require all three components working together.
@@ 5. Conclusion
We studied optimization failures in generative recommendation, focusing on the ”Symmetric Conservatism” induced by GBPO-style bounded updates, which empirically manifests as cold-start suppression and diversity collapse. To address this in list-wise generation, we proposed SAGE (Sequence-level Adaptive Gradient Evolution), featuring (i) Asymmetric Adaptive Dynamics via a Positive Boost for high-potential successful slates and an Entropy-Aware Penalty for low-diversity failures, and (ii) Multi-Objective Signal Decoupling to preserve gradient fidelity under mixed feedback.
Across three Amazon Product Reviews datasets and the large-scale industrial RecIF-Bench benchmark, TextRec—which reuses the LLM’s native vocabulary and is optimized with SAGE—consistently achieves the strongest overall accuracy while substantially improving long-tail coverage and recommendation diversity relative to OneRec-GBPO. The results also reveal a practical stability pattern: applying strong exploration to a constrained Semantic-ID action space can increase diversity but may sacrifice top-$K$ accuracy, whereas native-text generation better maintains the accuracy–diversity balance under SAGE.
A remaining limitation is that native-vocabulary recommendation requires language-level item rendering, which can be challenging for latency-sensitive systems; nevertheless, the consistent gains on both academic and industrial benchmarks suggest that native-text generation is a promising direction as compute and deployment tooling improve.